A survey of 51 pet owners revealed that 32 owned dogs, 38 owned cats, and 3 owned neither dogs nor cats. How many owned both dogs and cats?
Answer by edjones(8007)

51-3=48 own cars or dogs or both. Make a Venn diagram and put 3 in the rectangle but out side the circles.
Let x= owned both dogs and cats. That is the intersection of the 2 circles.
32-x+38-x+x=48 the union of the 2 circles IE everything in the circles.
32-x+38-x+x=48
70-x=48
-x=-22
x=22 owned both dogs and cats.
FYI: 10 goes to the left of the intersection and 16 to the right.
.
Ed